OpenStudy (anonymous):

Simplify quantity 2 times x plus 10 all over 2 2x + 5 x + 10 x + 5 x

OpenStudy (anonymous):

did u still need this 1?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

2x + 5 is my solution although see if someone can check it.

OpenStudy (anonymous):

@rmrjr22

OpenStudy (rmrjr22):

2(x+5)/2 = X+5....?

OpenStudy (anonymous):

@ganeshie8 brother can u check this please

ganeshie8 (ganeshie8):

2 times x plus 10 all over 2

ganeshie8 (ganeshie8):

lets break it

ganeshie8 (ganeshie8):

2 times x | plus 10 | all over 2

ganeshie8 (ganeshie8):

\(\large \frac{2x + 10}{2}\)

OpenStudy (rmrjr22):

tis the same thing

OpenStudy (anonymous):

so how would u simplify that down to?

ganeshie8 (ganeshie8):

divide 2 from each term in numerator term

OpenStudy (anonymous):

can u divide the 2x by 2?

OpenStudy (anonymous):

so its x + 5

OpenStudy (rmrjr22):

bingo!

OpenStudy (rmrjr22):

aww i said it first ... :(

ganeshie8 (ganeshie8):

ahh yes.... :)